# A Regret Perspective on Online Selective Generation

### v1

@@ F.3 Analysis on Time Horizon and Desired FDR Parameters
Figure 22: The violin plots for the ablation study over varying $\alpha$ on datasets under distribution-shift environments. The violin plots are drawn with randomly chosen dataset with single distribution shift between $60\mathrm{K}$-sized TriviaQA and $60\mathrm{K}$-sized NQ, over 30 random trials. Here, converged FDRs go along with target risk $\alpha$ (i.e., violin plots are below of the diagonal line), which empirically demonstrates the our algorithm finds an efficient selective generators with FDR guarantees. Note that $\textbf{Err}_{T}$ refers to the error rate on a given dataset, i.e., $\textbf{Err}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{1}\left(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)/T$ and the FDRs of selective generators should be lower than $\textbf{Err}_{T}$ to be meaningful in controlling the FDRs.
@@
Figure 23: Additional $\textbf{FDR}_{t}$ comparisons of selective generation methods under stochastic environments with a longer time horizon. Here, we repeat the datasets to evaluate whether Exp3-SG converges over a long horizon ($T=1,500\mathrm{K}$). As shown in the plots, Exp3-SG exhibits bare convergence after a prolonged horizon $T$.
Figure 24: Additional $\textbf{FDR}_{t}$ comparison of selective generation methods under distribution-shift environments with a longer time horizon. We consider a single distribution shift from $2000\mathrm{K}$-sized TriviaQA to $2000\mathrm{K}$-sized NQ, denoted in a dotted vertical line, along with GPT-3.5-turbo as our base generator ($T=4000\mathrm{K},\alpha=0.1$). As can be seen, Exp3-SG eventually satisfies a FDR even under the shift.
F.4 Ablation Study
@@
Figure 25: Comparison of ExSUL and ExSUL without the inefficiency loss $a_{t}$ (6) for the ablation study, with GPT-3.5-turbo as a generator on TriviaQA ($T=30\mathrm{K},\alpha=0.25$). Removing $a_{t}$ in ExSUL clearly leads to convergence to more inefficient solution. This behavior contrasts with the faster convergence of EW-SG compared to ExSUL or Exp3-SG , e.g., EW-SG and ExSUL eventually converges to similar levels of inefficiency in Figure 3. However, as seen in this plot, ExSUL without $a_{t}$ visibly converges to a much more inefficient generator, which demonstrates our algorithm provides an FDR guarantee while maximizing selection efficiency. Note that $\ell_{\text{max}}$ is set to $\lambda$, and the learning rate is also adjusted accordingly.
Appendix G Proofs
G.1 A Proof of Theorem 3
We follow the standard proof [13], followed by extending it for a loss range with $[0,\ell_{\text{max}}]$.
Let the cumulative loss of the $i$-th hypothesis $\tau_{i}\in\mathcal{{H}}$ be $L_{t}^{i}\coloneqq\sum_{s=1}^{t}\ell_{t}(\tau_{i})$ and $W_{t}\coloneqq\sum_{i=1}^{|\mathcal{{H}}|}\exp\left\{-\eta L_{t}^{i}\right\}$.
@@ G.1 A Proof of Theorem 3
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\leq\frac{\eta\ell_{\text{max}}^{2}T}{8}+\frac{\ln|\mathcal{{H}}|}{\eta}.$
Since the above upper bound of the regret is convex in $\eta$, we can obtain the optimal value of $\eta=\frac{2}{\ell_{\text{max}}}\sqrt{\frac{2\ln|\mathcal{{H}}|}{T}}$ that minimizes the upper bound, which completes the proof.
G.2 A Proof of Theorem 1
@@
This proof follows the standard proof [13] but with a non-trivial loss range $[0,\ell_{\text{max}}]$.
Recall that $\ell_{t}(\tau)\in[0,\ell_{\text{max}}]$ and its unbiased estimator ~$\tilde{\ell}_{t}(\tau\mid\tau_{t})\in[0,\infty)$, as defined in (2). Additionally, we define
$W_{t}\coloneqq\sum_{i=1}^{|\mathcal{{H}}|}\text{exp}(-\eta\tilde{L}_{t}^{i})\quad\text{where}\quad\tilde{\ell}_{t}^{i}\coloneqq\tilde{\ell}_{t}(\tau_{i}\mid\tau_{t})\text{~{}and~{}}\tilde{L}_{t}^{i}\coloneqq\sum_{s=1}^{t}\tilde{\ell}_{s}^{i}.$
The upper bound of potentials.
We find an upper bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. To this end, we first consider an upper bound of $\mathbbm{E}\ln\frac{W_{t}}{W_{t-1}}$.
$\ln\frac{W_{t}}{W_{t-1}}$ $=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t}^{i})}{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})}$
$=\ln\frac{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})\,\exp(-\eta\tilde{\ell}_{t}^{i})}{\sum_{i=1}^{|\mathcal{H}|}\exp(-\eta\,\tilde{L}_{t-1}^{i})}$
$=\ln\sum_{i=1}^{|\mathcal{H}|}\frac{\exp(-\eta\tilde{L}_{t-1}^{i})}{\sum_{j=1}^{|\mathcal{H}|}\exp(-\eta\tilde{L}_{t-1}^{j})}\exp(-\eta\tilde{\ell}_{t}^{i})$
$=\ln{\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})}\exp(-\eta\tilde{\ell}_{t}^{i})$ (11)
$\leq\ln\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(1-\eta\tilde{\ell}_{t}^{i}+\frac{\eta^{2}(\tilde{\ell}_{t}^{i})^{2}}{2}\right)$ (12)
$=\ln\left(1-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}\right)$
$\leq-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})(\tilde{\ell}_{t}^{i})^{2}$ (13)
@@
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\left(\frac{\ell_{t}^{i}}{p_{t}(\tau_{i})}\cdot\mathbbm{1}(\tau_{t}=\tau_{i})\right)\tilde{\ell}_{t}^{i}$ (14)
$=-\eta\sum_{i=1}^{|\mathcal{H}|}p_{t}(\tau_{i})\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{t}^{t}\tilde{\ell}_{t}^{t}$
$=-\eta\mathbbm{E}_{\tau_{i}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{t}^{t}\tilde{\ell}_{t}^{t}.$ (15)
Here, (11) follows directly from the definition of the sampling distribution $p_{t}(\tau)$. (12) is obtained by applying the second-order Taylor bound $e^{x}\leq 1+x+{x^{2}}/{2}$ which is valid for $x\leq 0$, together with the observation that $-\eta\tilde{\ell}_{t}^{i}\leq 0$. (13) holds by the standard logarithmic bound $\ln(1+x)\leq x$ for any $x\in\mathbb{R}$. Finally, (14) follows from the definition of our estimator.
Then, we have the following:
$\mathbbm{E}\left[\ln\frac{W_{T}}{W_{0}}\right]$ $=\mathbbm{E}\left[\sum_{t=1}^{T}\ln\frac{W_{t}}{W_{t-1}}\right]$
$\leq-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}^{t}\tilde{\ell}_{t}^{t}$ (16)
@@
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau_{t}\in\mathcal{{H}}}p_{t}(\tau_{t})\ell_{t}(\tau_{t})\left(\frac{\ell_{t}(\tau_{t})}{p_{t}(\tau_{t})}\mathbbm{1}(\tau_{t}=\tau_{t})\right)$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau\in\mathcal{{H}}}p_{t}(\tau)\ell_{t}(\tau)\frac{\ell_{t}(\tau)}{p_{t}(\tau)}$
$=-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\sum_{t=1}^{T}\sum_{\tau\in\mathcal{{H}}}\ell_{t}(\tau)^{2}$
$\leq-\eta\sum_{t=1}^{T}\mathbbm{E}_{\tau_{i},\tau_{t}\sim p_{t}}\tilde{\ell}_{t}^{i}+\frac{\eta^{2}}{2}\ell_{\text{max}}^{2}T|\mathcal{{H}}|$ (17)
After the telescoping sum, (16) applies the exponential-weights bound obtained in the previous step to each factor $\ln\frac{W_{t}}{W_{t-1}}$. Finally, (17) holds as $\ell_{t}(\cdot)$ is bounded by $\ell_{\text{max}}$.
The lower bound of potentials.
Then, we find a lower bound of $\mathbbm{E}\ln\frac{W_{T}}{W_{0}}$. In particular, we have

### v2

@@ G.3 Analysis on Time Horizon and Desired FDR Parameters
Figure 23: The violin plots for the ablation study over varying $\alpha$ on datasets under distribution-shift environments. The violin plots are drawn with randomly chosen dataset with single distribution shift between $15\mathrm{K}$-sized TriviaQA and $15\mathrm{K}$-sized NQ, over 30 random trials. Here, converged FDRs go along with target risk $\alpha$ (i.e., violin plots are below of the diagonal line), which empirically demonstrates the our algorithm finds an efficient selective generators with FDR guarantees. Note that $\textbf{Err}_{T}$ refers to the error rate on a given dataset, i.e., $\textbf{Err}_{T}\coloneqq\sum_{t=1}^{T}\mathbbm{1}\left(G(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t}\right)/T$ and the FDRs of selective generators should be lower than $\textbf{Err}_{T}$ to be meaningful in controlling the FDRs.
@@
Figure 24: Additional $\textbf{FDR}_{t}$ comparisons of selective generation methods under stochastic environments with a longer time horizon. Here, we repeat the datasets to evaluate whether Exp3-IX-SG converges over a long horizon ($T=1,500\mathrm{K}$). As shown in the plots, Exp3-IX-SG exhibits bare convergence after a prolonged horizon $T$.
(a) NQ dataset shift to TriviaQA dataset along with GPT-3.5-Turbo
(b) NQ dataset shift to TriviaQA dataset along with LLaMA3.1-8B
(c) TriviaQA dataset shift to NQ dataset along with GPT-3.5-Turbo
(d) TriviaQA dataset shift to NQ dataset along with LLaMA3.1-8B
Figure 25: Additional $\textbf{FDR}_{t}$ comparison of selective generation methods under distribution-shift environments with a longer time horizon. Here, we repeat the datasets to evaluate whether Exp3-IX-SG converges over a long horizon. We consider a single distribution shift between two $1000\mathrm{K}$-sized datasets (NQ, triviaQA) denoted in a dotted vertical line. As shown in the plots, Exp3-IX-SG exhibits bare convergence after a prolonged horizon $T$.
G.4 Analysis on Varying $\lambda$
(a) NQ dataset along with GPT-3.5-Turbo
(b) NQ dataset along with LLaMA3.1-8B
(c) TriviaQA dataset along with GPT-3.5-Turbo
(d) TriviaQA dataset along with LLaMA3.1-8B
Figure 26: Additional $\textbf{FDR}_{t}$ comparisons of selective generation methods under stochastic environments with varying $\lambda$ for $T=30\mathrm{K}$ datasets.
(a) NQ dataset along with GPT-3.5-Turbo
(b) NQ dataset along with LLaMA3.1-8B
(c) TriviaQA dataset along with GPT-3.5-Turbo
(d) TriviaQA dataset along with LLaMA3.1-8B
Figure 27: Additional $\textbf{Ineff}_{t}$ comparisons of selective generation methods under stochastic environments with varying $\lambda$ for $T=30\mathrm{K}$ datasets.
G.5 Ablation Study
@@
Figure 28: Comparison of ExSUL and ExSUL without the inefficiency loss $a_{t}$ (4) for the ablation study, with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},\alpha=0.25$). Removing $a_{t}$ in ExSUL clearly leads to convergence to more inefficient solution. This behavior contrasts with the faster convergence of EW-SG in Figure 8, compared to ExSUL or Exp3-SG , where EW-SG and ExSUL eventually converges to similar levels of inefficiency. As seen in the plot, ExSUL without $a_{t}$ visibly converges to a much more inefficient generator, demonstrating that our algorithm provides an FDR guarantee while maximizing selection efficiency.
Appendix H Proofs
H.1 A Proof of Theorem 3
We follow the standard proof (Foster & Rakhlin, 2023), followed by extending it for a loss range with $0,\ell_{\text{max}}]$.
Let the cumulative loss of the $i$-th hypothesis $\tau_{i}\in\mathcal{{H}}$ be $L_{t}^{i}\coloneqq\sum_{s=1}^{t}\ell_{t}(\tau_{i})$, $W_{t}\coloneqq\sum_{i=1}^{|\mathcal{{H}}|}\exp\left\{-\eta L_{t}^{i}\right\}$.
@@ H.1 A Proof of Theorem 3
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\leq\frac{\eta\ell_{\text{max}}^{2}T}{8}+\frac{\ln|\mathcal{{H}}|}{\eta}.$
If $\eta=\sqrt{8\ln|\mathcal{{H}}|/\ell_{\text{max}}^{2}T}$,
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)$ $\leq\ell_{\text{max}}\sqrt{\frac{T\ln|\mathcal{{H}}|}{2}}.$ (13)
Since the guarantee in (13) holds for any sequence of loss functions in EW, the bound is also valid against an adaptive adversary. Furthermore, the expected regret is given by $\mathbbm{E}\left[\textbf{Reg}_{T}\right]=\mathbbm{E}\left[\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\right],$ which follows from the tower property of expectation.
High probability bound. Moreover, we have
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)$ $=\left(\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\right)+\left(\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\right).$
Since $\mathbbm{E}_{t}\left[\ell_{t}(\tau_{t})-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\right]=0$, which implies that $\sum_{t=1}^{T}\left[\ell_{t}(\tau_{t})-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\right]$ is a martingale, we can use Hoeffding-Azuma inequality
$\mathbbm{P}\left(\sum_{t=1}^{T}\left[\ell_{t}(\tau_{t})-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\right]\geq\varepsilon\right)\leq\exp\left(-\frac{2\varepsilon^{2}}{\sum_{t=1}^{T}c_{t}^{2}}\right)$
for all $\varepsilon\geq 0$ where $c_{t}=\ell_{\text{max}}$ since $\ell_{t}(\tau_{t})-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\in[-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t}),\ell_{\text{max}}-\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})]$.
Thus, if we put $\delta=\exp\left(-\frac{2\varepsilon^{2}}{T\ell_{\text{max}}^{2}}\right)$ we get
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})\leq\ell_{\text{max}}\sqrt{\frac{T\ln\delta^{-1}}{2}},$
with probability at least $1-\delta$, which suggests that
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)$ $\leq\frac{\eta\ell_{\text{max}}^{2}T}{8}+\frac{\ln|\mathcal{{H}}|}{\eta}+\ell_{\text{max}}\sqrt{\frac{T\ln\delta^{-1}}{2}}$
with probability at least $1-\delta$.
If $\eta=\sqrt{8\ln|\mathcal{{H}}|/\ell_{\text{max}}^{2}T}$,
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)$ $\leq\ell_{\text{max}}\sqrt{\frac{T\ln|\mathcal{{H}}|}{2}}+\ell_{\text{max}}\sqrt{\frac{T\ln\delta^{-1}}{2}}.$ (14)
H.2 A Proof of Theorem 2
@@
This proof follows standard proof (Neu, 2015) but with a non-trivial loss range $0,\ell_{\text{max}}]$. Recall that $\ell_{t}(\tau)\in[0,\ell_{\text{max}}]$ and its biased estimator $\tilde{\ell}_{t}(\tau\mid\{\tau_{t}\})\in[0,\infty)$, as defined in (9).
We first introduce new Lemma 2 and Corollary 23 that establish a high probability bound between the loss estimator and the empirical loss. Corollary 23 provides a bound for a fixed value of $\tau$, and its proof follows directly from the Lemma 2 by applying a union bound. Together, Lemma 2 and Corollary 23 allow us to translate the upper bound on the estimator equation into a high-probability regret bound.
Lemma 2.
Let $\gamma_{t}\geq 0$ for $t\in\{1,2,\dots,T\}$ be a fixed non-increasing sequence and $\alpha_{t,\tau}$ be a real value that satisfies $\alpha_{t,\tau}\leq 2\gamma_{t}$ for all $t\in\mathbb{N}$ and $\tau\in\mathcal{{H}}$. Then, we have
$\sum_{t=1}^{T}\sum_{\tau\in\mathcal{{H}}}\alpha_{t,\tau}\left(\ell_{t}(\tau\mid\{\tau_{t}\})-\ell_{t}(\tau)\right)\leq\ell_{\text{max}}\ln\frac{1}{\delta},$
with probability $1-\delta$.
Proof.
First, letting $\beta_{t}=\frac{2\gamma_{t}}{\ell_{\text{max}}}$, we have
$\ell_{t}(\tau\mid\{\tau_{t}\})$ $=\frac{\ell_{t}(\tau)}{p_{t}(\tau)+\gamma_{t}}\cdot\mathbbm{1}(\tau_{t}=\tau)$
$=\ell_{\text{max}}\cdot\frac{\ell_{t}(\tau)/\ell_{\text{max}}}{p_{t}(\tau)+\gamma_{t}}\cdot\mathbbm{1}(\tau_{t}=\tau)$
$\leq\ell_{\text{max}}\cdot\frac{\ell_{t}(\tau)/\ell_{\text{max}}}{p_{t}(\tau)+\gamma_{t}\cdot\ell_{t}(\tau)/\ell_{\text{max}}}\cdot\mathbbm{1}(\tau_{t}=\tau)$
$=\frac{\ell_{\text{max}}}{2\gamma_{t}}\cdot\frac{2\gamma_{t}\cdot\frac{\ell_{t}(\tau)}{\ell_{\text{max}}p_{t}(\tau)}}{1+\gamma_{t}\cdot\frac{\ell_{t}(\tau)}{\ell_{\text{max}}p_{t}(\tau)}}\mathbbm{1}(\tau_{t}=\tau)$
$\leq\frac{\ell_{\text{max}}}{2\gamma_{t}}\cdot\ln\left(1+2\gamma_{t}\cdot\frac{\ell_{t}(\tau)}{\ell_{\text{max}}p_{t}(\tau)}\mathbbm{1}(\tau_{t}=\tau)\right)$ (15)
$=\frac{1}{\beta_{t}}\cdot\ln\left(1+\beta_{t}\frac{\ell_{t}(\tau)}{p_{t}(\tau)}\mathbbm{1}(\tau_{t}=\tau)\right)$ (16)
where the inequality (15) holds by $\frac{x}{1+x/2}\leq\ln(1+x)$ for $x\geq 0$ and the range $\{0,1\}$ of the indicator function.
Second, we explicitly define the following auxiliary quantities:
$\tilde{S}_{t}=\sum_{\tau\in\mathcal{{H}}}\frac{\alpha_{t,\tau}}{\ell_{\text{max}}}\ell_{t}(\tau\mid\{\tau_{t}\})\quad\text{and}\quad S_{t}=\sum_{\tau\in\mathcal{{H}}}\frac{\alpha_{t,\tau}}{\ell_{\text{max}}}\ell_{t}(\tau).$
Denoting by $\mathbbm{E}_{t}$ the conditional expectation with respect to $\tau_{t}$, given the past learner’s choices $\tau_{1},\ldots,\tau_{t-1}$, we have the following relation between $\tilde{S}_{t}$ and $S_{t}$:
@@
$\mathbbm{E}_{t}\left[\exp\tilde{S}_{t}\right]$ $\leq\mathbbm{E}_{t}\left[\exp\left(\sum_{\tau\in\mathcal{{H}}}\frac{\alpha_{t,\tau}}{\ell_{\text{max}}\beta_{t}}\cdot\ln\left(1+\beta_{t}\cdot\frac{\ell_{t}(\tau)}{p_{t}(\tau)}\mathbbm{1}(\tau_{t}=\tau)\right)\right)\right]$ (17)
@@
$=\mathbbm{E}_{t}\left[\prod_{\tau\in\mathcal{{H}}}\left(1+\beta_{t}\frac{\ell_{t}(\tau)}{p_{t}(\tau)}\mathbbm{1}(\tau_{t}=\tau)\right)^{\frac{\alpha_{t,\tau}}{\ell_{\text{max}}\beta_{t}}}\right]$